# Prompt cusps in hierarchical dark matter halos: Implications for annihilation boost

### v1

@@ Abstract
Recent simulations have identified long-lived “prompt cusps”—compact remnants of early density peaks with inner profiles $\rho\propto r^{-3/2}$. They can survive hierarchical assembly and potentially enhance signals of dark matter annihilation. In this work, we incorporate prompt cusps into the semi-analytic substructure framework SASHIMI, enabling a fully hierarchical, environment-dependent calculation of the annihilation luminosity that consistently tracks subhalos, sub-subhalos, and tidal stripping. We assign prompt cusps to first-generation microhalos and propagate their survival through the merger history, including an explicit treatment of cusps associated with stripped substructure. We find that the substructure hierarchy converges rapidly once a few levels are included, and that prompt cusps can raise the total annihilation boost of Milky-Way–size hosts at $z=0$ to $B\sim\mathcal{O}(10)$ for fiducial cusp-occupation assumptions, compared to a subhalo-only baseline of $B_{\rm sh}\sim\mathrm{few}$. Across a wide range of host masses and redshifts, prompt cusps increase the normalization of $B(M_{\rm host},z)$ while largely preserving its mass and redshift trends. Compared to universal-average, peak-based estimates, our fiducial boosts are lower by about an order of magnitude, primarily reflecting a correspondingly smaller inferred cusp abundance in host halos, highlighting the importance of unifying peak-based cusp formation with merger-tree evolution and environmental dependence.
@@ 1 Introduction
Numerical simulations have long indicated that CDM halos follow a nearly universal density profile, commonly parametrized by the Navarro-Frenk-White (NFW) form, $\rho(r)\propto r^{-1}$ in the inner regions and $\rho(r)\propto r^{-3}$ at large radii [32]. However, recent high-resolution simulations have revealed a qualitatively different behavior in the earliest-forming halos. Small-scale peaks in the initial density field can undergo rapid, nearly self-similar collapse, producing extremely steep inner cusps with $\rho(r)\propto r^{-1.5}$—significantly steeper than the NFW profile [26, 1, 27, 15, 17, 18]. These prompt cusps appear generically in the first generation of halos, with their number and properties determined by the statistics of primordial density peaks. Only peaks with sufficiently low ellipticity and without strong interactions with neighboring structures form such cusps; others relax to an NFW-like structure instead. Importantly, simulations suggest that once formed, prompt cusps tend to survive subsequent mergers and accretion events, remaining as long-lived features embedded in the inner regions of larger halos [15].
The existence of prompt cusps has significant implications for indirect searches for dark matter. Because the annihilation rate scales with the square of the local dark matter density, the extremely steep $r^{-1.5}$ profiles generated by the rapid collapse of small-scale peaks can substantially enhance the gamma-ray emissivity associated with WIMP annihilation. Delos and White [18] demonstrated that, for thermally produced WIMPs, every solar mass of dark matter may contain thousands of Earth-mass prompt cusps, leading to an overall annihilation rate that exceeds predictions based on smooth-halo or traditional subhalo models by at least an order of magnitude. Their analysis further showed that prompt-cusp emission is less centrally concentrated than the conventional NFW-based expectation, thereby modifying the relative importance of diffuse components such as the extragalactic gamma-ray background [18]. Subsequent works test these predictions against observations. Constraints from the isotropic gamma-ray background measured by the Fermi-LAT have been shown to strongly limit the viable parameter space for WIMP annihilation once prompt cusps are included [17]. Additional studies have considered environmental effects such as disruption by stars and tidal fields, finding that cusp survival can significantly influence the expected annihilation signal [42]. Observations of galaxy clusters have been used to place further constraints on prompt-cusp annihilation scenarios, in some cases suggesting substantial tension between cusp-enhanced predictions and current gamma-ray limits [13].
@@ 3.2 Hierarchical substructure and effective prompt-cusp counts
In hierarchical structure formation, dark matter (sub)halos can be associated, to a good approximation, with peaks of the underlying (filtered) Gaussian density field, so that each subhalo realization may be linked to a corresponding peak in the initial conditions [5, 11, 12]. Motivated by this peak-based picture, we assume that a prompt cusp can form around a peak only if the peak collapses sufficiently early, and that only a fraction $f_{\rm coll}$ of peaks satisfies this collapse condition (as discussed in the previous subsection). Moreover, not all newly formed prompt cusps are expected to survive subsequent evolution: tidal processing and mergers can erase a non-negligible fraction of cusps, leading to an overall survival probability of $f_{\rm surv}\simeq 0.4$$0.6$, as suggested by cosmological considerations of cusp survival [18].
Combining these ingredients, we parametrize the fraction of subhalos that host a surviving prompt cusp as
$f_{\rm cusp}\equiv f_{\rm coll}\,f_{\rm surv},$ (3.7)
such that $f_{\rm cusp}$ may be directly compared to the effective cusp fraction adopted in Ref. [18]. For reference, our peak-collapse estimate gives $f_{\rm coll}\simeq 0.48$, and together with $f_{\rm surv}\simeq 0.4$$0.6$ this suggests $f_{\rm cusp}\sim 0.2$$0.3$, making $f_{\rm cusp}=0.25$ a plausible benchmark value. In what follows, we therefore explore $f_{\rm cusp}=\{1,\ 0.5,\ 0.25\}$ as representative choices spanning optimistic to conservative scenarios for cusp survival within the subhalo population.
@@ 3.3 Tidal stripping of hierarchical substructure and cusp survival
We therefore split the effective number of $\mathrm{sub}^{n}$-subhalos into two components, $N_{\rm sh,in}^{(n)}$ (inside $r_{t}$ of their immediate parent, hence retained as bound members of the hierarchy) and $N_{\rm sh,stripped}^{(n)}$ (initially outside $r_{t}$, hence stripped). In our fiducial implementation, stripped subhalos are assumed to be disrupted and phase-mixed into the host component, so that their associated cusps do not contribute as identifiable substructures. To allow for the possibility that a fraction of cusps in stripped material could remain self-bound and survive as compact remnants, we introduce an additional survival parameter $f_{\rm surv,stripped}$, defined as the fraction of prompt cusps associated with stripped subhalos that persist after stripping This ingredient is complementary to existing treatments that focus on the tidal stripping of the cusps themselves and on cusp disruption/merging criteria. For example, Ref. [18] explicitly models tidal stripping of prompt cusps and discuss cusp survival relative to subhalo disruption, but do not track the fate of a hierarchical sub-subhalo population split by a truncation radius within a semi-analytic subhalo catalog.
For a Milky-Way–size host ($M=10^{12}M_{\odot}$, $z=0$), the resulting effective subhalo counts are summarized in Table 1. The number of surviving cusps are then obtained by
$N_{\rm cusp}=f_{\rm cusp}\left(N_{\rm sh,in}+f_{\rm surv,stripped}N_{\rm sh,stripped}\right).$ (3.9)
To bracket the (currently uncertain) resilience of cusps associated with $\mathrm{sub}^{n}$-subhalos that are released when their parent branch is tidally stripped, we explore representative choices for this parameter. The case $f_{\rm surv,stripped}=0$ corresponds to the most conservative limit, where cusps tied to substructure in the stripped region are assumed to be fully destroyed or efficiently mixed into the smooth host component. On the other hand, $f_{\rm surv,stripped}=1$ is the maximally optimistic limit in which cusps survive stripping essentially unscathed, so that all cusp contributions are retained regardless of whether their parent subhalos remain bound.
@@ 4 Results
Figure 1: Mass-normalized annihilation luminosity parameters at $z=0$ including hierarchical substructure up to $\mathrm{sub}^{4}$-subhalos, shown as functions of the host-halo mass $M_{\rm host}$. The dot-dashed, dashed, and solid curves show the contributions from the smooth host-halo component, the total subhalo population, and the total prompt-cusp population, respectively (i.e., $J_{\rm host}/M_{\rm host}$, $J_{\rm sh,,total}/M_{\rm host}$, and $J_{\rm cusp,,total}/M_{\rm host}$), for the fiducial choice of $f_{\rm cusp}=0.25$ and $f_{\rm surv,stripped}=1$. The dotted curve shows the intrinsic cusp annihilation efficiency, $\langle J_{\rm cusp}\rangle/\langle M_{\rm cusp}\rangle$.
Figure 1 shows the mass-normalized luminosity parameters that enter the boost-factor calculation, plotted as functions of the host mass at $z=0$. The host contribution $J_{\rm host}/M_{\rm host}$ decreases with increasing $M_{\rm host}$, reflecting the systematic reduction in halo concentration (and hence in $\rho^{2}$-weighted mass) toward higher masses. The subhalo contribution $J_{\rm sh,\,total}/M_{\rm host}$ varies more gently with $M$, as it represents an integral over the subhalo population whose internal structure and abundance evolve only mildly across the mass range considered. In contrast, the prompt-cusp contribution $J_{\rm cusp,\,total}/M_{\rm host}$ remains comparatively large over a broad range of host masses for our fiducial parameters ($f_{\rm cusp}=0.25$, $f_{\rm surv,stripped}=1$), illustrating that even a moderate surviving cusp fraction can yield a sizeable contribution to the total annihilation budget.
@@
Figure 2 summarizes our main result for the annihilation boost factor at $z=0$ when hierarchical substructure is included up to $\mathrm{sub}^{4}$-subhalos. The colored bands visualize two independent modeling ingredients for prompt cusps. First, the overall vertical placement of each band is controlled by the cusp-occupation fraction $f_{\rm cusp}$, i.e., the fraction of (sub)halos that host a surviving cusp. Second, the band width captures the additional uncertainty associated with cusps that originate from subhalos located outside the tidal truncation radius and are subsequently stripped: we bracket this effect by varying $f_{\rm surv,stripped}\in[0,1]$, where $f_{\rm surv,stripped}=0$ assumes complete destruction of stripped-region cusps, while $f_{\rm surv,stripped}=1$ assumes they survive and continue to contribute. For comparison, the dotted curve shows the standard subhalo-only prediction without prompt cusps, illustrating that prompt cusps can enhance the boost but remain modest for fiducial parameter choices.
Figure 3 illustrates how the predicted boost changes as progressively deeper levels of hierarchical substructure are included. The sequence from $n=0$ to $n=4$ shows that the hierarchy rapidly approaches a converged result once a few substructure levels are accounted for; i.e. the incremental impact of adding ever higher-order sub-substructure becomes small. For comparison, the dotted curve shows the no-cusp baseline evaluated at $n=4$.
Figure 4 shows the redshift evolution of the boost factor. At fixed host mass, the boost varies with $z$ due to the changing halo structural properties and the shorter time available for hierarchical buildup and tidal processing at earlier epochs. In all cases, the prompt-cusp contribution (solid) shifts the boost above the no-cusp baseline (dotted), while preserving the overall mass trend. This figure therefore provides a compact summary of how both the subhalo hierarchy and prompt cusps jointly shape $B(M_{\rm host},z)$.
@@ 5.1 Comparison with earlier work
For similar benchmark assumptions as in the universal-average treatment of Ref. [18] ($f_{\rm cusp}=0.25$ and $f_{\rm surv,stripped}=1$), our predicted Milky-Way–size boost is $B\sim\mathcal{O}(10)$ at $z=0$, about one order of magnitude smaller than the $\sim 100$ quoted by Delos and White [18]. This offset can be understood primarily as a difference in the inferred abundance of surviving cusps: their Fig. 4 suggests $N_{\rm cusp}\sim 4\times 10^{16}$ in a $10^{12}M_{\odot}$ halo, whereas Table 1 combined with Eq. (3.9) yields $N_{\rm cusp}\sim 3\times 10^{15}$. Since $J_{\rm cusp,total}\propto N_{\rm cusp}\langle J_{\rm cusp}\rangle$ at fixed cusp internal structure, a one-dex difference in $N_{\rm cusp}$ maps directly onto a one-dex difference in $B$. Equivalently, adopting the typical prompt-cusp mass $\langle M_{\rm cusp}\rangle\sim 10^{-6}M_{\odot}$ (Eq. 2.15 and surrounding discussion), the implied total mass bound in cusps is $M_{\rm cusp,tot}\sim N_{\rm cusp}\langle M_{\rm cusp}\rangle\sim 4\times 10^{10}M_{\odot}$ in the Delos and White estimate (i.e. $\sim 4\%$ of a $10^{12}M_{\odot}$ host), while our hierarchical SASHIMI-based count implies $M_{\rm cusp,tot}\sim 3\times 10^{9}M_{\odot}$ (i.e. $\sim 0.3\%$ of the host mass).
@@ 5.2 Possible origin of the cusp-abundance difference
Second, Ref. [18] effectively counts cusps associated with peaks that collapse down to $z=0$. In contrast, our implementation assigns cusps at the formation of (sub)halos within the hierarchical tree and does not create additional cusps within an already-accreted branch; subsequent evolution can only reduce the identifiable cusp population through stripping or disruption. If late-time cusp formation in already-accreted material is non-negligible, our procedure will undercount cusps relative to a universal peak-based estimate.
Third, microhalo-scale structures that host primordial cusps are expected to be weakly clustered ($b<1$) relative to dark matter particles in many environments, whereas a universal-average mapping effectively assumes an unbiased distribution when translating “cusps per unit mass” into “cusps per host” independent of environment. Even a modest anti-bias could reduce the expected cusp abundance in Milky-Way–like regions relative to the universal estimate.
Lastly, an additional, and potentially important, source of uncertainty in the EPS-based microhalo abundance concerns the mapping between the cutoff scale and halo mass when computing $\sigma(M)$ for a suppressed small-scale power spectrum. In SASHIMI we evaluate $\sigma(M)$ with a sharp-$k$ filter and adopt a cutoff $k_{\max}=\alpha/R$, where $\alpha$ parametrizes the residual ambiguity of this mapping [38]. Our fiducial choice, $\alpha=1.8$, is motivated by matching to the real-space top-hat result in the large-mass regime, but it is not calibrated directly on microhalo simulations. Crucially, varying $\alpha$ can change the predicted abundance of microhaloes—and hence the inferred number of prompt cusps—without dramatically affecting the mass function on the high-mass end that dominates typical calibrations. For example, we find that choosing $\alpha=1$ increases the predicted cusp abundance by a factor of $\sim 5$ relative to the fiducial model, while $\alpha=2.5$ reduces it to $\sim 0.4$ times the fiducial value This sensitivity suggests that a dedicated calibration of $\alpha$ in the microhalo regime (e.g. using ultra-high-resolution simulations that resolve Earth-mass haloes and their environmental dependence) will be important for robust predictions of $N_{\rm cusp}$ and, in turn, of the cusp-induced annihilation boost.

### v2

@@ Abstract
Recent simulations have identified long-lived “prompt cusps”—compact remnants of early density peaks with inner profiles $\rho\propto r^{-3/2}$. They can survive hierarchical assembly and potentially enhance signals of dark matter annihilation. In this work, we incorporate prompt cusps into the semi-analytic substructure framework SASHIMI, enabling a fully hierarchical, environment-dependent calculation of the annihilation luminosity that consistently tracks subhalos, sub-subhalos, and tidal stripping. We assign prompt cusps to first-generation microhalos and propagate their survival through the merger history, including an explicit treatment of cusps associated with stripped substructure. We find that the substructure hierarchy converges rapidly once a few levels are included, and that prompt cusps can raise the total annihilation boost of Milky-Way–size hosts at $z=0$ to $B\sim 50$ for fiducial cusp-occupation assumptions, compared to a subhalo-only baseline of $B_{\rm sh}\sim\mathrm{few}$. Across a wide range of host masses and redshifts, prompt cusps increase the normalization of $B(M_{\rm host},z)$ while largely preserving its mass and redshift trends. Compared to universal-average, peak-based estimates, our fiducial boosts are lower by about a factor of a few, primarily reflecting a correspondingly smaller inferred cusp abundance in host halos, highlighting the importance of unifying peak-based cusp formation with merger-tree evolution and environmental dependence.
@@ 1 Introduction
Numerical simulations have long indicated that CDM halos follow a nearly universal density profile, commonly parametrized by the Navarro-Frenk-White (NFW) form, $\rho(r)\propto r^{-1}$ in the inner regions and $\rho(r)\propto r^{-3}$ at large radii [33]. However, recent high-resolution simulations have revealed a qualitatively different behavior in the earliest-forming halos. Small-scale peaks in the initial density field can undergo rapid, nearly self-similar collapse, producing extremely steep inner cusps with $\rho(r)\propto r^{-1.5}$—significantly steeper than the NFW profile [27, 1, 28, 15, 35, 17, 18, 21]. These prompt cusps appear generically in the first generation of halos, with their number and properties determined by the statistics of primordial density peaks. Only peaks with sufficiently low ellipticity and without strong interactions with neighboring structures form such cusps; others relax to an NFW-like structure instead. Importantly, simulations suggest that once formed, prompt cusps tend to survive subsequent mergers and accretion events, remaining as long-lived features embedded in the inner regions of larger halos [15].
The existence of prompt cusps has significant implications for indirect searches for dark matter. Because the annihilation rate scales with the square of the local dark matter density, the extremely steep $r^{-1.5}$ profiles generated by the rapid collapse of small-scale peaks can substantially enhance the gamma-ray emissivity associated with WIMP annihilation. Delos and White [18] demonstrated that, for thermally produced WIMPs, every solar mass of dark matter may contain thousands of Earth-mass prompt cusps, leading to an overall annihilation rate that exceeds predictions based on smooth-halo or traditional subhalo models by at least an order of magnitude. Their analysis further showed that prompt-cusp emission is less centrally concentrated than the conventional NFW-based expectation, thereby modifying the relative importance of diffuse components such as the extragalactic gamma-ray background [18]. Subsequent works test these predictions against observations. Constraints from the isotropic gamma-ray background measured by the Fermi-LAT have been shown to strongly limit the viable parameter space for WIMP annihilation once prompt cusps are included [17]. Additional studies have considered environmental effects such as disruption by stars and tidal fields, finding that cusp survival can significantly influence the expected annihilation signal [45]. Observations of galaxy clusters have been used to place further constraints on prompt-cusp annihilation scenarios, in some cases suggesting substantial tension between cusp-enhanced predictions and current gamma-ray limits [13, 34].
@@ 3.2 Hierarchical substructure and effective prompt-cusp counts
In hierarchical structure formation, dark matter halos can be associated, to a good approximation, with peaks of the underlying Gaussian density field [5, 11, 12]. However, not every peak in the initial density field gives rise to a bound halo. Only peaks that collapse sufficiently early and reach high central densities form self-bound halos, and it is these collapsed peaks that can host a central prompt cusp. Cosmological arguments and numerical studies indicate that roughly half of the initial peaks satisfy this collapse condition, corresponding to a collapse fraction $f_{\rm coll}\simeq 0.48$ (see Sec. 2.4), while the remainder never form bound halos and therefore do not contribute to the subhalo population. In this work, we therefore adopt the physically motivated assumption that every collapsed primordial halo contains a central prompt cusp.
We introduce $f_{\rm cusp}$ as a dimensionless parameter that denotes the fraction of halos whose central prompt cusps contribute effectively to the annihilation signal. Our fiducial choice is $f_{\rm cusp}=1$, corresponding to the assumption that all collapsed halos host cusps that remain relevant for annihilation. We retain $f_{\rm cusp}$ as a free parameter to allow for the possibility that, in a fraction of halos, prompt cusps become sufficiently modified by hierarchical evolution that their annihilation contribution is strongly suppressed. Exploring $f_{\rm cusp}<1$ therefore provides a conservative way to bracket uncertainties in the impact of cusp evolution on the annihilation signal.
@@ 3.3 Tidal stripping of hierarchical substructure and cusp survival
We therefore split the effective number of $\mathrm{sub}^{n}$-subhalos into two components: $N_{\rm sh,in}^{(n)}$, corresponding to subhalos located inside the truncation radius $r_{t}$ of their immediate parent and thus retained as bound members of the hierarchy, and $N_{\rm sh,stripped}^{(n)}$, corresponding to subhalos initially outside $r_{t}$ that are released by tidal stripping. Physically, the central cusps associated with stripped subhalos are expected to be resilient against tidal forces. However, once released into the host halo, the extent to which these cusps remain identifiable as independent contributors to the annihilation signal is uncertain within a halo-based semi-analytic description. To bracket this uncertainty, we introduce an effective parameter $f_{\rm surv,stripped}$, defined as the fraction of prompt cusps associated with stripped subhalos that contribute independently to the annihilation signal.
For a Milky-Way–size host ($M=10^{12}M_{\odot}$, $z=0$), the resulting effective subhalo counts are summarized in Table 1. The number of cusps contributing to the annihilation signal is then written as
$N_{\rm cusp}=f_{\rm cusp}\left(N_{\rm sh,in}+f_{\rm surv,stripped}\,N_{\rm sh,stripped}\right),$ (3.8)
where, as discussed in the previous subsection, $f_{\rm cusp}$ controls the overall fraction of halos whose central cusps contribute effectively to the annihilation signal. Physically, cusps are expected to survive tidal stripping. Accordingly, we adopt $f_{\rm surv,stripped}=1$ as the fiducial choice throughout this work. We nevertheless explore values $f_{\rm surv,stripped}<1$ as a conservative way to allow for the possibility that cusps associated with stripped substructure contribute less efficiently to the annihilation signal, for example due to orbital heating, etc.
@@ 4 Results
Figure 1: Mass-normalized annihilation luminosity parameters at $z=0$ including hierarchical substructure up to $\mathrm{sub}^{4}$-subhalos, shown as functions of the host-halo mass $M_{\rm host}$. The dot-dashed, dashed, and solid curves show the contributions from the smooth host-halo component, the total subhalo population, and the total prompt-cusp population, respectively (i.e., $J_{\rm host}/M_{\rm host}$, $J_{\rm sh,,total}/M_{\rm host}$, and $J_{\rm cusp,,total}/M_{\rm host}$), for the fiducial choice of $f_{\rm cusp}=f_{\rm surv,stripped}=1$. The dotted curve shows the intrinsic cusp annihilation efficiency, $\langle J_{\rm cusp}\rangle/\langle M_{\rm cusp}\rangle$.
Figure 1 shows the mass-normalized luminosity parameters that enter the boost-factor calculation, plotted as functions of the host mass at $z=0$. The host contribution $J_{\rm host}/M_{\rm host}$ decreases with increasing $M_{\rm host}$, reflecting the systematic reduction in halo concentration (and hence in $\rho^{2}$-weighted mass) toward higher masses. The subhalo contribution $J_{\rm sh,\,total}/M_{\rm host}$ varies more gently with $M_{\rm host}$, as it represents an integral over the subhalo population whose internal structure and abundance evolve only mildly across the mass range considered. In contrast, the prompt-cusp contribution $J_{\rm cusp,\,total}/M_{\rm host}$ remains comparatively large over a broad range of host masses for our fiducial parameters ($f_{\rm cusp}=f_{\rm surv,stripped}=1$), illustrating that the cusps can yield a dominant contribution to the total annihilation budget.
@@
Figure 2 shows the resulting annihilation boost factor at $z=0$ for the converged hierarchical model, adopting the fiducial parameter choice $f_{\rm cusp}=1$ and $f_{\rm surv,stripped}=1$. Relative to the standard subhalo-only prediction (dotted curve), the inclusion of prompt cusps enhances the boost while remaining modest for Milky-Way–size hosts. It also illustrates how the predicted annihilation boost changes as progressively deeper levels of hierarchical substructure are included. The sequence from $n=0$ to $n=4$ shows that the boost rapidly approaches a converged result once a few substructure levels are accounted for, indicating that the hierarchical expansion is well controlled. This figure therefore summarizes the main physical impact of prompt cusps in a fully hierarchical setting.
Figure 3 shows the redshift evolution of the boost factor for the same fiducial model. At fixed host mass, the boost varies with redshift due to the evolving structural properties of halos and the reduced time available for hierarchical buildup and tidal processing at earlier epochs. In all cases, the inclusion of prompt cusps shifts the boost above the no-cusp baseline while preserving the overall mass dependence.
@@
Finally, we assess the sensitivity of our results to the modelling assumptions associated with prompt cusps. In Fig. 4, the colored bands indicate the range obtained by varying the cusp-occupation fraction $f_{\rm cusp}$ and the stripped-cusp contribution parameter $f_{\rm surv,stripped}$ around their fiducial values. The overall normalization of the boost is controlled by $f_{\rm cusp}$, while the band width reflects the additional uncertainty associated with cusps originating from stripped substructure.
@@ 5.1 Comparison with earlier work
For similar benchmark assumptions as in the universal-average treatment of Ref. [18] ($f_{\rm cusp}=f_{\rm surv,stripped}=1$), our predicted Milky-Way–size boost is $B\sim 50$ at $z=0$, about a factor of four smaller than the $\sim 200$ quoted by Delos and White [18]. This offset can be understood primarily as a difference in the inferred abundance of surviving cusps: they have $N_{\rm cusp}\sim 2\times 10^{16}$ in a $10^{12}M_{\odot}$ halo, whereas Table 1 combined with Eq. (3.8) yields $N_{\rm cusp}\sim 10^{16}$. Since $J_{\rm cusp,total}\propto N_{\rm cusp}\langle J_{\rm cusp}\rangle$ at fixed cusp internal structure, this difference in $N_{\rm cusp}$ maps directly onto that in $B$. Equivalently, adopting the typical prompt-cusp mass $\langle M_{\rm cusp}\rangle\sim 10^{-6}M_{\odot}$ (Eq. 2.15 and surrounding discussion), the implied total mass bound in cusps is $M_{\rm cusp,tot}\sim N_{\rm cusp}\langle M_{\rm cusp}\rangle\sim 2\times 10^{10}M_{\odot}$ in the Delos and White estimate (i.e. $\sim 2\%$ of a $10^{12}M_{\odot}$ host), while our hierarchical SASHIMI-based count implies $M_{\rm cusp,tot}\sim 10^{10}M_{\odot}$ (i.e. $\sim 1\%$ of the host mass).
@@ 5.2 Possible origin of the cusp-abundance difference
Second, microhalo-scale structures that host primordial cusps are expected to be weakly clustered ($b<1$) relative to dark matter particles in many environments, whereas a universal-average mapping effectively assumes an unbiased distribution when translating “cusps per unit mass” into “cusps per host” independent of environment. Even a modest anti-bias could reduce the expected cusp abundance in Milky-Way–like regions relative to the universal estimate.
Lastly, an additional, and potentially important, source of uncertainty in the EPS-based microhalo abundance concerns the mapping between the cutoff scale and halo mass when computing $\sigma(M)$ for a suppressed small-scale power spectrum. In SASHIMI we evaluate $\sigma(M)$ with a sharp-$k$ filter and adopt a cutoff $k_{\max}=\alpha/R$, where $\alpha$ parametrizes the residual ambiguity of this mapping [41]. Our fiducial choice, $\alpha=1.8$, is motivated by matching to the real-space top-hat result in the large-mass regime, but it is not calibrated directly on microhalo simulations. Crucially, varying $\alpha$ can change the predicted abundance of microhaloes—and hence the inferred number of prompt cusps—without dramatically affecting the mass function on the high-mass end that dominates typical calibrations. For example, we find that choosing $\alpha=1$ increases the predicted cusp abundance by a factor of $\sim 5$ relative to the fiducial model, while $\alpha=2.5$ reduces it to $\sim 0.4$ times the fiducial value. This sensitivity suggests that a dedicated calibration of $\alpha$ in the microhalo regime (e.g. using ultra-high-resolution simulations that resolve Earth-mass haloes and their environmental dependence) will be important for robust predictions of $N_{\rm cusp}$ and, in turn, of the cusp-induced annihilation boost.